### v3

# ClinKD: Cross-Modal Clinical Knowledge Distiller For Multi-Task Medical Images

Abstract
Medical Visual Question Answering (Med-VQA) represents a critical and challenging subtask within the general VQA domain. Despite significant advancements in general Visual Question Answering (VQA), multimodal large language models (MLLMs) still exhibit substantial limitations when handling multi-task VQA scenarios. These limitations manifest through erroneous spatial localization and misinterpretation of medical images, which primarily arise from two fundamental issues: inadequate image-text alignment and insufficient medical knowledge in general-purpose MLLMs for specialized medical applications. To address these issues, we introduce the Cross-Modal Clinical Knowledge Distiller (ClinKD), an innovative framework designed to enhance image-text alignment and establish more effective medical knowledge adaptation mechanisms, which enables MLLMs to adapt to medical knowledge. Our extensive experimental evaluations demonstrate that the ClinKD achieves state-of-the-art performance on the Med-GRIT-270k dataset, a challenging medical benchmark containing fine-grained multi-task QA pairs. The results indicate that our approach not only significantly improves image-text alignment but also effectively enables MLLMs to adapt to the medical knowledge. The source code for ClinKD is available at: https://github.com/overloadedHenry/ClinKD. Refer to caption Figure 1: The left part shows that the BiRD [14] only utilizes traditional supervised fine-tuning (SFT) and 2D-RoPE [32] to obtain the capabilities of understanding and grounding medical images. The right part is our method which uses Med-CLIP Guided RoPE for enhancing image-text alignment and Pseudo-Labels Medical Knowledge Distillation for filling the gap of medical knowledge in ClinKD. During the SFT procedure, we will use Reflective Correction Training to enable answers with low semantic similarity or incorrect groundings to be reused.
1 Introduction
Medical Visual Question Answering (Med-VQA) [25, 1, 10] is a critical domain for applying MLLMs [37, 21, 4] to medical image analysis [19]. Recent advancements in MLLMs have enabled them to achieve preliminary capabilities in image analysis, annotation, and user instruction compliance [5]. Previous studies have made multiple attempts [38, 15, 11, 29, 18, 14] to enhance the performance of general-purpose MLLMs in Med-VQA.
For instance, LLaVA-Med [18] extends instruction-tuning techniques for MLLMs to the medical domain. This work applied instruction tuning and curated image-text datasets to align visual-textual features in LLaVA [27], enabling LLaVA-Med to recognize and analyze medical images. However, this system lacked fine-grained image understanding, causing erroneous spatial localization when questioned about specific pathological locations [9, 14].
Subsequently, the BiRD [14], a model that incorporated fine-grained image recognition by developing a biomedical refer-and-ground instruction-tuning dataset, addresses the absence of grounding capabilities [40, 42] in LLaVA-Med’s work [18]. Nevertheless, as shown in Figure 1, due to its reliance on traditional supervised fine-tuning (instruction-tuning) [28, 17, 45] and conventional 2D Rotary Position Embedding (RoPE) [32], BiRD tends to generate erroneous spatial localization and misinterpretation of medical images because of incomplete image-text alignment [35, 43]. Moreover, the visual encoder of BiRD is constrained by the lack of medical knowledge, resulting in difficulties in medical knowledge adaptation [47, 24].
To overcome these limitations, we propose a medical distillation framework named ClinKD to boost capabilities in the Med-VQA domain. Specifically, our method mainly focuses on two issues: incomplete image-text alignment and the professional medical knowledge gap between the general-purpose MLLMs and medical domain that requires finer-grained medical knowledge [22, 41].
For the first issue, we propose Med-CLIP Guided Rotary Position Embedding (MCG-RoPE), a novel position embedding method that focuses on the inter-modal and intra-modal intervals during joint image-text training, while ensuring the equivalence of context intervals immediately adjacent to images. Compared to traditional RoPE [32], the MCG-RoPE uses the distinct index intervals to better capture cross-modal information, thereby achieving more effective image-text alignment.
For the second issue, we propose Clinical Knowledge Distiller (ClinKD) which consists of Pseudo-Labels Medical Knowledge Distillation and Reflective Correction Training. Pseudo-Labels Medical Knowledge Distillation provides prior medical knowledge by using pseudo-labels [20, 6], filling the gap of prior medical knowledge so that the model can better adapt to medical knowledge during the supervised fine-tuning. Reflective Correction Training allows samples with low semantic similarity to be sent for training again after being enhanced by MLLMs.
Our contributions are summarized as follows:
  • We introduce the ClinKD, an innovative distiller designed to establish more effective medical knowledge adaptation mechanisms.
  • For solving incomplete image-text alignment, we propose the Med-CLIP Guided Rotary Position Embedding (MCG-RoPE). The MCG-RoPE utilizes the distinct inter-modal and intra-modal features to improve image-text alignment.
  • For bridging medical knowledge gap between general-purpose MLLMs and specialized medical applications, we propose the Pseudo-Labels Medical Knowledge Distillation, filling the medical knowledge gap before supervised fine-tuning.
  • Extensive experimental results show that the proposed ClinKD achieves the best performance on Med-GRIT-270k [14].
Figure 2: Overview of our ClinKD’s framework: The framework begins with multi-task medical images being segmented into patches. The med-extractor module with Med-CLIP Guided Rotary Position Embedding extracts dimensional features from the patched image embeddings. The ClinKD system undergoes distillation before supervised fine-tuning (SFT). During the SFT, answers with low semantic similarity or incorrect groundings are prioritized for retraining. When responding to queries, the framework selects the answer with the highest score, ensuring accurate medical image analysis and interpretation.
2 Related Work
2.1 Biomedical Instruction-Tuning Datasets
Many Med-VQA datasets such as VQA-RAD [16] and SLAKE [26] contain only medical images and simple QA pairs. Specifically, these QA pairs do not have the fine-grained descriptions of pathological features. This limitation constrains the ability of MLLMs to perform accurate spatial localization. To boost and evaluate MLLMs’ capabilities in the Med-VQA domain, many biomedical instruction-tuning datasets are designed.
For example, the LLaVA-Med-qa0.2k [46] contains 50 medical images sampled from several modalities such as CT, Angiography and PET and 193 questions according to these medical images. The medical QA pairs are generated by GPT-4 based on the metadata. This dataset is designed to evaluate the capabilities of describing medical images, specifically the biomedical features.
Unlike the LLaVA-Med-qa0.2k [46], the Med-GRIT-270k [14] includes multi-task fine-grained medical QA pairs especially for grounding spatial locations. It is designed to enhance the model’s capabilities in referring and grounding. This multi-task dataset comprises four distinct tasks: Visual Grounding (VG), Referring Object Classification (ROC), Referring Captioning (RC), and Medical Image Analysis (MIA). The VG task evaluates the model’s capability of accurately matching text descriptions to corresponding image regions. The RC task assesses the model’s capability to recognize specific image areas and generate descriptive captions for them. The ROC task examines the model’s understanding of textual information related to particular image regions and their associated visual details. The MIA task evaluates the model’s comprehension of medical images and their multi-modal context. All data is sourced from eight modalities: CT, MR, X-ray, PET, Endoscopy, Dermoscopy, Fundus, and Ultrasound. Utilizing data from these eight modalities allows for a more comprehensive evaluation of the model’s capabilities.
2.2 Multimodal Large Language Models (MLLMs) for Biomedicine
With the development of MLLMs, an increasing number of approaches have been applied to the medical field, such as BioMedGPT [44], LLaVA-Med [18], and BiRD [14]. These methods drive the development of MLLMs in the biomedical domain from different perspectives. For instance, LLaVA-Med [18] applied instruction fine-tuning to the medical imaging field and built a medical question-answering system based on LLaVA [27], which provided reasonable responses by processing user-submitted images and instructions. However, LLaVA-Med did not support fine-grained interactions with medical images, meaning that the system cannot precisely locate pathological regions within the images. To address the lack of referring and grounding capabilities, the BiRD [14], which is based on Qwen-VL [3], leveraged a finer-grained medical dataset. However, the ViT of BiRD was not trained by sufficient medical data before being frozen, resulting in a gap in medical knowledge domain. Compared with their work [14, 18], our method modifies the position embedding method and introduces an effective pseudo-labels medical knowledge distillation framework.
2.3 Rotary Position Embedding (RoPE)
The introduced RoPE [32] makes the relative position features of the context in the text more easily captured by LLMs. Further, the RoPE-Mixed [12] introduced mixed learnable frequencies to RoPE, enabling it to handle diagonal directions and making RoPE itself learnable. It showed more efficiency when dealing with diagonal features of the image. The VideoRoPE [36] extended the diagonal layout and variable frequency to the video domain, demonstrating its advantages in video analysis tasks. Unlike existing methods, our approach focuses on 2D multi-task medical images where capturing semantic information is challenging. We utilize cross-modal intervals to enable the model to better distinguish modal information, thereby improving image-text alignment.
3 Methodology
Figure 3: The visualized example of Med-CLIP Guided Rotary Position Embedding. The original vector will be rotated by different angles according to the transformed position indexes. The inter-modal and intra-modal intervals are different so that the Med-CLIP Guided Rotary Position Embedding can better align image feature and text feature.
As shown in Figure 2, we aim at enhancing the ClinKD’s performance by mainly leveraging Med-CLIP Guided Rotary Position Embedding, Pseudo-Labels Medical Knowledge Distillation, Reflective Correction Training and Semantic-Aware Selective Generation. The proposed Med-CLIP Guided Rotary Position uses the distinct inter-modal and intra-modal space to improve image-text alignment. The distiller provides prior medical knowledge to enhance the model’s capability of medical knowledge adaptation. In the inference procedure, the Semantic-Aware Selective Generation will allow the model to choose the answer with the best score.
3.1 Med-CLIP Guided Rotary Position Embedding
The 2D Rotary Position Embedding (RoPE) [32] tends to lack attention on cross-modal features, so we propose the Med-CLIP Guided Rotary Position Embedding (MCG-RoPE) to modify rotation angles that are crucial for image-text alignment. Figure 3 shows the example of our method.
Distinct Inter-Modal Intervals. Suppose we have a sentence-image-sentence position index sequence:
$S_{idx}=($ $T_{1},T_{2}\cdots,T_{\ell},V_{\ell+1},V_{\ell+2},\cdots,$ (1)
$V_{\ell+wh},T_{\ell+wh+1},T_{\ell+wh+2},\cdots,T_{\ell+wh+k})$
where $T_{i}=(t_{i},t_{i})$ denotes the 2D position index of text tokens, $V_{i}=(v^{(1)}_{i},v^{(2)}_{i})$ denotes the 2D image index.
To distinguish between intra-modal features and inter-modal features, the indexes of image tokens need to be transformed.
$\left(v_{i}^{(1)},v_{i}^{(2)}\right)\xrightarrow{\mathcal{F}}\left(\alpha_{1}v_{i}^{(1)}+\lambda_{1},\alpha_{2}v_{i}^{(2)}+\lambda_{2}\right)=\mathcal{V}_{i}$ (2)
where $\mathcal{V}_{i}$ is the position after being scaled and shifted. The $\lambda_{1},\lambda_{2},\alpha_{1},\alpha_{2}$ are parameters. The $\mathcal{F}$ denotes the operator of shifting and scaling.
The inter-modal intervals should be consistent, and thus the following equation needs to be satisfied:
$\begin{cases}\alpha_{1}+\lambda_{1}-\ell=\ell+wh+1-(\alpha_{1}h+\lambda_{1})\\ \alpha_{2}+\lambda_{2}-\ell=\ell+wh+1-(\alpha_{2}w+\lambda_{2})\end{cases}$ (3)
We need to fix $\alpha_{1},\alpha_{2}$ first to set the intervals between each image’s position index. So the general solution is
$\begin{cases}\lambda_{1}=\ell+\frac{[wh+1-\alpha_{1}(h+1)]}{2}\\ \lambda_{2}=\ell+\frac{[wh+1-\alpha_{2}(w+1)]}{2}\end{cases}$ (4)
After setting intervals, the new sequence can be expressed as:
$S_{h-idx}=($ $t_{1},\cdots,t_{\ell},\alpha_{1}v^{(1)}_{\ell+1}+\lambda_{1},\cdots,$ (5)
$\alpha_{1}v^{(1)}_{\ell+wh}+\lambda_{1},,t_{\ell+wh+1},\cdots,t_{\ell+wh+k})$
where $S_{h-idx}$ is the position index on height dimension,
$S_{w-idx}=($ $t_{1},\cdots,t_{\ell},\alpha_{2}v^{(2)}_{\ell+1}+\lambda_{2},\cdots,$ (6)
$\alpha_{2}v^{(2)}_{\ell+wh}+\lambda_{2},,t_{\ell+wh+1},\cdots,t_{\ell+wh+k})$
where $S_{w-idx}$ is the position index on width dimension.
Distinct Interval-Based Rotation. Suppose we have a query/key vector: $q\in\mathbb{R}^{d_{h}+d_{w}}$ where $\mathbb{R}^{d_{h}+d_{w}}$ denotes the $d_{h}+d_{w}$ dimension Euclidean space. The $d_{h}$, $d_{w}$ denotes the embedding dimension on height dimension and width dimension respectively. Then $q$ will be split into $q^{(h)}\in\mathbb{R}^{d_{h}}$ and $q^{(w)}\in\mathbb{R}^{d_{w}}$
The corresponding frequency on the two dimensions is set as:
$\omega_{i}^{(h)}=10000^{-\frac{2i}{d_{h}}}$ (7)
$\omega_{i}^{(w)}=10000^{-\frac{2i}{d_{w}}}$
Then the angles of rotation are expressed as:
$\theta_{x_{h},i}^{(h)}=x_{h}\cdot\omega_{i}^{(h)}$ (8)
$\theta_{x_{w},i}^{(w)}=x_{w}\cdot\omega_{i}^{(w)}$
where $x_{h}$, $x_{w}$ denote the transformed position index on height and width dimension respectively. This means the angle of rotation will be determined by inter-modal and intra-modal intervals. A large rotation angle brings information of modal transformation while a smaller angle contains fine-grained features of intra-modal. The rotation matrix can be defined as:
$\boldsymbol{R}_{\boldsymbol{\Theta},x_{h}}^{d_{h}}=\begin{pmatrix}R_{\theta_{x_{h},1}^{(h)}}&&\\ &\ddots&\\ &&R_{\theta_{x_{h},d_{h}/2}^{(h)}}\end{pmatrix}$ (9)
$\boldsymbol{R}_{\boldsymbol{\Theta},x_{w}}^{d_{w}}=\begin{pmatrix}R_{\theta_{x_{w},1}^{(w)}}&&\\ &\ddots&\\ &&R_{\theta_{x_{w},d_{w}/2}^{(w)}}\end{pmatrix}$
where $R_{\theta_{x_{h},i}^{(h)}}$ and $R_{\theta_{x_{w},i}^{(w)}}$ are defined as:
$R_{\theta_{x_{w},i}^{(w)}}=\begin{pmatrix}{\operatorname{cos}\theta_{x_{w},i}^{(w)}}&{-\operatorname{sin}\theta_{x_{w},i}^{(w)}}\\ {\operatorname{sin}\theta_{x_{w},i}^{(w)}}&{\operatorname{cos}\theta_{x_{w},i}^{(w)}}\\ \end{pmatrix}$ (10)
$R_{\theta_{x_{h},i}^{(h)}}=\begin{pmatrix}{\operatorname{cos}\theta_{x_{h},i}^{(h)}}&{-\operatorname{sin}\theta_{x_{h},i}^{(h)}}\\ {\operatorname{sin}\theta_{x_{h},i}^{(h)}}&{\operatorname{cos}\theta_{x_{h},i}^{(h)}}\\ \end{pmatrix}$
The rotated vector is shown in Eq. (11) :
$\widetilde{q}=\left[\boldsymbol{R}_{\boldsymbol{\Theta},x_{h}}^{d_{h}}{q^{(h)}};\boldsymbol{R}_{\boldsymbol{\Theta},x_{h}}^{d_{h}}{q^{(w)}}\right]$ (11)
3.2 Pseudo-Labels Medical Knowledge Distillation (Pseudo-KD)
Our clinical knowledge distiller enables the student model to learn the output distribution of the teacher model. Therefore, we provide prior knowledge of medical images to the student model through Pseudo-Labels Medical Knowledge Distillation.
The distillation loss function is defined by a linear combination of cross-entropy loss and the Kullback-Leibler Divergence (KL Divergence).
The cross-entropy loss is defined as:
$\mathcal{L}_{CE}=-\sum_{i=1}^{N}y_{i}\log p_{s}(z_{s_{i}})$ (12)
where $\mathcal{L}_{CE}$ denotes the cross-entropy loss, $p_{s}(z_{i})$ is the logits generated by the student ClinKD model and $y_{i}$ is the hard label of ground truth.
The KL Divergence mentioned before is defined as:
$\mathcal{L}_{KD}=T^{2}\mathbb{D}^{\boldsymbol{\epsilon_{t}},\boldsymbol{\epsilon_{s}}}_{KL}(p_{t}\|p_{s})$ (13)
where $\mathcal{L}_{KD}$ denotes the loss function constructed by KL Divergence, $\mathbb{D}^{\boldsymbol{\epsilon_{t}},\boldsymbol{\epsilon_{s}}}_{KL}(p_{t}\|p_{s})$ is the KL Divergence, $T$ is called temperature which can smooth the distribution and the $\epsilon_{s_{j}}$, $\epsilon_{t_{j}}$ are noise values.
Furthermore, the $\mathbb{D}^{\boldsymbol{\epsilon_{t}},\boldsymbol{\epsilon_{s}}}_{KL}(p_{t}\|p_{s})$ is:
$\mathbb{D}^{\boldsymbol{\epsilon_{t}},\boldsymbol{\epsilon_{s}}}_{KL}(p_{t}\|p_{s})=\sum_{i=1}^{n}p_{t}(z_{t_{i}}+\epsilon_{t_{j}})\log\frac{p_{t}(z_{t_{i}}+\epsilon_{t_{j}})}{p_{s}(z_{s_{i}}+\epsilon_{s_{j}})}$ (14)
where $p_{t}(z_{i})$ is the probability distribution generated by the teacher ClinKD model, $p_{s}(z_{i})$ denotes the probability distribution generated by the student ClinKD model.
For details, the distributions are smoothed by temperature $T$:
$p_{s}(z_{s_{j}}+\epsilon_{s_{j}})=\dfrac{exp((z_{s_{j}}+\epsilon_{s_{j}})/T)}{\displaystyle\sum_{i}^{n}exp((z_{s_{i}}+\epsilon_{s_{i}})/T)}$ (15)
$p_{t}(z_{t_{j}}+\epsilon_{t_{j}})=\dfrac{exp((z_{t_{j}}+\epsilon_{t_{j}})/T)}{\displaystyle\sum_{i}^{n}exp((z_{t_{i}}+\epsilon_{t_{i}})/T)}$
Finally the whole distillation loss is expressed as:
$\mathcal{L}_{Distill}=\dfrac{\mathcal{L}_{CE}+\alpha\mathcal{L}_{KD}}{1+\alpha}$ (16)
After obtaining teacher’s prior knowledge about medical images, the student ClinKD model will be trained by supervised fine-tuning (SFT).
3.3 Reflective Correction Training
After the model distillation, we restart two rounds of supervised fine-tuning using the whole dataset. During the supervised fine-tuning process, the model’s outputs are recorded and compared with the ground truth. If the semantic similarity between the output and the ground truth falls below 80%, GPT-4o will correct the output based on the ground truth (retaining only the original meaning and anchor box coordinates) to enrich the diversity of the training set labels. The corrected data is then fed back into the training process for further learning.
3.4 Semantic-Aware Selective Generation
The responses generated by large language models exhibit randomness, and relying on a single inference may result in missing the optimal answer. Semantic-Aware Selective Generation means that instead of directly using the model’s single-round generation for evaluation, we allow the model to generate multiple samples. We then utilize the CLIP models to score the alignment between the generated text and image. Finally, we reorder the outputs based on these scores and select the one with the highest score as the final result. The pseudo code is shown in Algorithm 1.
Algorithm 1 Semantic-Aware Selective Generation
1:  Input: Image $I$, Model $M$
2:  Output: Optimal generated text $T_{final}$
3:  $T_{1},T_{2},\dots,T_{n}\leftarrow\text{GenerateSamples}(M,I)$
4:  for $i=1$ to $n$ do
5:     $S_{i}\leftarrow\text{CLIPScore}(T_{i},I)$
6:  end for
7:  $T_{ordered}\leftarrow\text{Sort}(T_{1},T_{2},\dots,T_{n})$
8:  $T_{final}\leftarrow T_{ordered}[1]$
9:  return  $T_{final}$
4 Experiments
Model Test dataset VG (Recall@0.5) ROC (Recall) RC (SPICE) MIA (mBMR) Average
LLaVA-Med [18] Med-GRIT-Test30k [39, 7] 0 2.75 8.18 11.20 5.53
BiRD [14] Med-GRIT-Test30k [39, 7] 53.92 65.33 55.23 52.17 56.66
ClinKD (Ours) Med-GRIT-Test30k [39, 7] 67.51 82.35 70.56 65.69 71.53
LLaVA-Med [18] LLaVA-Med-qa0.2k [46] - - - 20.04 -
BiRD [14] LLaVA-Med-qa0.2k [46] - - - 10.55 -
ClinKD (Ours) LLaVA-Med-qa0.2k [46] - - - 13.24 -
Table 1: Comparison with LLaVA-Med [18] and BiRD (previous SOTA) [14]. The best scores are shown in bold.
Model Proposed Methods VG (Recall@0.5) ROC (Recall) RC (SPICE) MIA (mBMR) Average
MCG-RoPE Pseudo-KD SASG
Qwen2-VL [34] 55.43 (+1.51) 68.33 (+3.00) 57.39 (+2.16) 62.23 (+10.06) 61.37 (+4.71)
BiRD [14] 53.92 65.33 55.23 52.17 56.66
64.03 (+10.11) 72.74 (+7.41) 61.32 (+6.09) 59.63 (+7.46) 64.44 (+7.78)
66.33 (+12.41) 79.52 (+14.19) 67.21 (+11.98) 64.32 (+12.15) 69.35 (+12.69)
55.89 (+1.97) 67.84 (+2.51) 54.32 (-0.91) 51.27 (-0.90) 57.33 (+0.67)
ClinKD (Ours) 67.51 (+13.59) 82.35 (+17.02) 70.56 (+15.33) 65.69 (+13.52) 71.53 (+14.87)
Table 2: Ablation study for different models and proposed methods. The value which is blue shows how much the metric scores improved. The best scores are shown in bold. The MCG-RoPE, Pseudo-KD and SASG stand for Med-CLIP Guided RoPE, Pseudo-Labels Medical Knowledge Distillation and Semantic-Aware Selective Generation respectively.
Task Metric Method CT MR X-ray PET Endoscopy Dermoscopy Fundus Ultrasound Average
BiRD [14] 44.47$\pm$0.14 29.26$\pm$0.11 41.73$\pm$0.16 56.46$\pm$0.15 53.60$\pm$0.12 75.63$\pm$0.17 84.15$\pm$0.13 46.04$\pm$0.18 53.92$\pm$0.14
BiRD [14] + RoPE-Mixed [12] 46.90$\pm$0.16 53.11$\pm$0.13 42.03$\pm$0.17 60.84$\pm$0.14 53.49$\pm$0.15 78.25$\pm$0.11 85.16$\pm$0.12 52.83$\pm$0.13 59.08$\pm$0.19
BiRD [14] + MCG-RoPE (Ours) 51.22$\pm$0.14 49.40$\pm$0.18 45.92$\pm$0.17 68.60$\pm$0.13 60.37$\pm$0.19 88.28$\pm$0.12 86.59$\pm$0.15 61.88$\pm$0.16 64.03$\pm$0.15
BiRD [14] + $V_{k}D$  [30] 58.77$\pm$0.12 50.95$\pm$0.16 45.69$\pm$0.11 65.89$\pm$0.18 57.63$\pm$0.14 93.32$\pm$0.15 79.27$\pm$0.12 59.90$\pm$0.15 63.93$\pm$0.19
BiRD [14] + Pseudo-KD (Ours) 62.59$\pm$0.15 52.86$\pm$0.17 47.19$\pm$0.13 66.28$\pm$0.16 59.84$\pm$0.11 93.68$\pm$0.18 85.37$\pm$0.14 62.87$\pm$0.19 66.33$\pm$0.13
VG Recall@0.5 ClinKD (Ours) 65.52$\pm$0.13 52.23$\pm$0.30 48.56$\pm$0.19 69.25$\pm$0.12 60.37$\pm$0.16 94.22$\pm$0.10 86.59$\pm$0.14 64.47$\pm$0.15 67.51$\pm$0.18
BiRD [14] 34.76$\pm$0.11 61.79$\pm$0.15 53.74$\pm$0.14 - 60.40$\pm$0.17 96.61$\pm$0.18 - 84.65$\pm$0.12 65.33$\pm$0.14
BiRD [14] + RoPE-Mixed [12] 39.78$\pm$0.13 62.13$\pm$0.19 74.27$\pm$0.11 - 61.33$\pm$0.13 96.61$\pm$0.17 - 84.65$\pm$0.15 69.80$\pm$0.12
BiRD [14] + MCG-RoPE (Ours) 44.47$\pm$0.15 64.20$\pm$0.18 83.75$\pm$0.11 - 62.19$\pm$0.14 96.61$\pm$0.19 - 85.22$\pm$0.16 72.74$\pm$0.13
BiRD [14] + $V_{k}D$ [30] 53.19$\pm$0.17 69.10$\pm$0.14 90.84$\pm$0.19 - 61.71$\pm$0.10 100$\pm$0.00 - 81.15$\pm$0.12 76.00$\pm$0.15
BiRD [14] + Pseudo-KD (Ours) 58.47$\pm$0.16 69.92$\pm$0.12 95.83$\pm$0.18 - 66.32$\pm$0.11 100$\pm$0.00 - 86.58$\pm$0.13 79.52$\pm$0.14
ROC Recall ClinKD (Ours) 61.42$\pm$0.17 71.30$\pm$0.19 97.00$\pm$0.18 - 68.84$\pm$0.12 100$\pm$0.00 - 95.54$\pm$0.11 82.35$\pm$0.15
BiRD [14] 41.88$\pm$0.12 51.69$\pm$0.14 37.39$\pm$0.18 47.95$\pm$0.15 54.07$\pm$0.10 77.44$\pm$0.16 48.73$\pm$0.13 82.65$\pm$0.19 55.23$\pm$0.14
BiRD [14] + RoPE-Mixed [12] 43.51$\pm$0.15 51.88$\pm$0.18 54.61$\pm$0.13 52.83$\pm$0.19 54.07$\pm$0.11 68.15$\pm$0.17 53.99$\pm$0.12 65.28$\pm$0.14 55.54$\pm$0.13
BiRD [14] + MCG-RoPE (Ours) 50.73$\pm$0.14 53.11$\pm$0.16 60.33$\pm$0.12 70.16$\pm$0.18 54.07$\pm$0.11 70.18$\pm$0.19 61.52$\pm$0.15 70.46$\pm$0.17 61.32$\pm$0.13
BiRD [14] + $V_{k}D$  [30] 51.33$\pm$0.13 54.39$\pm$0.17 62.83$\pm$0.19 74.31$\pm$0.14 54.07$\pm$0.15 76.23$\pm$0.12 67.89$\pm$0.18 72.15$\pm$0.11 64.15$\pm$0.16
BiRD [14] + Pseudo-KD (Ours) 52.78$\pm$0.12 58.41$\pm$0.15 70.39$\pm$0.18 75.85$\pm$0.13 55.22$\pm$0.16 81.34$\pm$0.11 71.54$\pm$0.17 72.15$\pm$0.14 67.21$\pm$0.19
RC SPICE [2] ClinKD (Ours) 53.94$\pm$0.14 58.41$\pm$0.12 71.47$\pm$0.18 79.91$\pm$0.15 52.10$\pm$0.11 83.92$\pm$0.19 84.51$\pm$0.13 80.19$\pm$0.16 70.56$\pm$0.17
BiRD [14] 47.01$\pm$0.14 49.35$\pm$0.18 37.17$\pm$0.15 57.15$\pm$0.12 39.91$\pm$0.17 72.13$\pm$0.13 48.87$\pm$0.16 65.78$\pm$0.19 52.17$\pm$0.11
BiRD [14] + RoPE-Mixed [12] 57.10$\pm$0.17 50.89$\pm$0.13 42.55$\pm$0.16 69.28$\pm$0.11 41.63$\pm$0.18 73.32$\pm$0.14 59.63$\pm$0.15 61.80$\pm$0.10 57.03$\pm$0.19
BiRD [14] + MCG-RoPE (Ours) 60.53$\pm$0.16 53.77$\pm$0.11 54.10$\pm$0.18 62.01$\pm$0.12 43.88$\pm$0.15 75.33$\pm$0.14 62.89$\pm$0.19 64.53$\pm$0.13 59.63$\pm$0.17
BiRD [14] + $V_{k}D$  [30] 62.33$\pm$0.10 53.77$\pm$0.15 52.27$\pm$0.19 62.01$\pm$0.16 47.29$\pm$0.13 75.33$\pm$0.17 68.04$\pm$0.12 70.31$\pm$0.18 61.42$\pm$0.14
BiRD [14] + Pseudo-KD (Ours) 66.85$\pm$0.11 55.73$\pm$0.14 52.27$\pm$0.17 64.74$\pm$0.15 50.58$\pm$0.12 78.21$\pm$0.19 70.02$\pm$0.13 76.16$\pm$0.18 64.32$\pm$0.16
MIA mBMR [14] ClinKD (Ours) 66.85$\pm$0.13 56.82$\pm$0.16 54.01$\pm$0.18 67.08$\pm$0.14 50.58$\pm$0.11 81.00$\pm$0.15 70.02 $\pm$0.17 76.05$\pm$0.19 65.69$\pm$0.12
BiRD [14] 43.03$\pm$0.12 48.02$\pm$0.11 42.51$\pm$0.17 53.85$\pm$0.19 51.99$\pm$0.14 80.45$\pm$0.16 60.58$\pm$0.13 69.78$\pm$0.15 -
BiRD [14] + RoPE-Mixed [12] 46.82$\pm$0.17 54.50$\pm$0.12 53.37$\pm$0.15 60.98$\pm$0.13 52.63$\pm$0.18 79.08$\pm$0.10 66.26$\pm$0.19 66.14$\pm$0.11 -
BiRD [14] + MCG-RoPE (Ours) 51.74$\pm$0.13 55.12$\pm$0.19 61.03$\pm$0.12 66.92$\pm$0.17 55.13$\pm$0.14 82.60$\pm$0.15 70.33$\pm$0.11 70.52$\pm$0.18 -
BiRD [14] + $V_{k}D$ [30] 56.41$\pm$0.16 57.05$\pm$0.11 62.91$\pm$0.19 67.40$\pm$0.14 55.18$\pm$0.18 86.22$\pm$0.10 71.73$\pm$0.15 70.88$\pm$0.13 -
BiRD [14] + Pseudo-KD (Ours) 60.17$\pm$0.14 59.23$\pm$0.16 66.42$\pm$0.12 68.96$\pm$0.19 57.99$\pm$0.13 88.31$\pm$0.15 75.64$\pm$0.17 74.44$\pm$0.10 -
Average - ClinKD (Ours) 61.93$\pm$0.15 59.69$\pm$0.18 67.76$\pm$0.19 72.08$\pm$0.13 57.97$\pm$0.16 89.79$\pm$0.11 80.37$\pm$0.14 79.06$\pm$0.12 -
Table 3: Comparison with other SOTA methods on Med-GRIT-Test30k [39, 7]. We evaluate the performance on eight modalities.
4.1 Datasets
Med-GRIT-270k [14]. The Med-GRIT-270k dataset comprises 270k samples, with 240k allocated for training and the remaining 30k for multi-task evaluation. These samples are collected from eight medical modalities: CT, MR, X-ray, PET, Endoscopy, Dermoscopy, Fundus and Ultrasound. Additionally, we randomly select 30K samples from the entire dataset for pseudo-label generation, which is subsequently used by the Pseudo-Labels Medical Knowledge Distillation. This fine-grained multi-task dataset not only demands that the model accurately understands medical images, but also requires it to precisely classify and ground pathological regions and physiological structures. This poses a significant challenge for MLLMs.
LLaVA-Med-qa0.2k [46]. The LLaVA-Med-qa0.2k [46] contains 0.2k QA pairs of MIA task which are used for evaluating the models’ ability to analyze medical images. It consists of 50 images with caption pairs and two types of questions: conversation and detailed description.
The images are not only from common modalities such as CT and MR, but also from some modalities that are not available in the Med-GRIT-270k [14] dataset, such as 3D reconstruction. The wide range of data sources evaluates the capabilities of MLLMs in general medical scenarios.
4.2 Evaluation Metrics
We use Recall@0.5 for Visual Grounding (VG), Recall for Referring Object Classification (ROC), SPICE [2] for Referring Caption (RC), and mBMR [14] for Medical Image Analysis (MIA). The final overall result is computed as the mean of the evaluation scores across these four tasks.
For the evaluation on LLaVA-Med-qa0.2k [18] which has 0.2k QA pairs of MIA task, we use mBMR as the metric.
4.3 Implementation Details
We use Qwen2-VL [34] as the base model for ClinKD. For the Med-CLIP Guided RoPE component of ClinKD, in order to ensure that the position index is an integer, we resample the image resolution to $w=h$. During the distillation process, we set $\alpha=0.5$, $T=0.5$. The number of distillation epochs is 2, with an additional 2 epochs for subsequent training. The entire training process is conducted using 6 NVIDIA A100 40G GPUs. The learning rate is set to $2\times 10^{-5}$, and the AdamW optimizer is employed. The learning scheduler follows a cosine schedule. For model generation, we use non-sampling beam search with a beam size of 3.
4.4 Analysis
4.4.1 Comparison with the State-of-the-Art Methods
Evaluation on the Med-GRIT-Test30k [39, 7]. As shown in Table 1, we compare ClinKD with BiRD [14] and LLaVA-Med [18]. Results show that ClinKD achieves scores of 67.51%, 82.35%, 70.56%, and 65.69% in the VG, ROC, RC, and MIA tasks respectively, outperforming other comparisons by a clear margin in all tasks. On average, ClinKD outperforms BiRD [14] and LLaVA-Med [18] by 14.87% and 66.00% respectively. ClinKD’s advantages lie in its more adequate image-text alignment and the prior medical knowledge provided by Pseudo-Labels Medical Knowledge Distillation.
Evaluation on the LLaVA-Med-qa0.2k [46]. Table 1 represents the models’ performance on LLaVA-Med-qa0.2k [46]. LLaVA-Med [18] still achieves the best performance. Although ClinKD has not yet surpassed LLaVA-Med [18] on this dataset, it still achieved a 2.69% improvement over BiRD [14] in the MIA task. The limitation may be caused by the different sources of medical data so that the ClinKD may fail to leverage medical knowledge to analyze medical images.
Figure 4: Effect of value of $\alpha$ in Pseudo-Labels Medical Knowledge Distillation.
4.4.2 Ablation study
Effect of Qwen2-VL [34]. Table 2 presents the ablation study of proposed methods. To investigate the impact of model upgrade on the experiment, we applied the same training strategy to Qwen2-VL as we did to BiRD [14]. The results show that the model upgrade only has a significant effect on the MIA task, which shows Qwen2-VL [34] outperforms BiRD by 10.06%.
Effect of Med-CLIP Guided Rotary Position Embedding. As shown in Table 2, with adding MCG-RoPE to BiRD, all metrics exhibit significant enhancements, with the average rising from 56.66% to 64.44%. This indicates the importance of adequate image-text alignment.
Effect of Pseudo-Labels Medical Knowledge Distillation. Table 2 represents improvements achieved by using Pseudo-Labels Medical Knowledge Distillation that leads to enhancements by 12.41%, 14.19%, 11.98% and 12.15% in the VG, ROC, RC and MIA tasks, respectively. The results illustrate that providing prior medical knowledge by using Pseudo-Labels Medical Knowledge Distillation can fill MLLMs’ gap of medical knowledge, boosting capabilities of adapting to new medical knowledge.
Figure 4 shows the effect of different values of $\alpha$ in Eq. (16). In this experiment, the average performance achieves the best when we set $\alpha=0.5$. Both excessively high and low alpha values can lead to performance degradation. This might be because a low $\alpha$ value prevents ClinKD from learning sufficient prior medical knowledge, while an overly high alpha value tends to make ClinKD generate irrational pseudo-labels that contradict common medical knowledge.
Effect of Semantic-Aware Selective Generation. From the results in Table 2, we can see a light enhancement on VG and ROC task. However, there is a slight decrease in the metric scores on the RC and MIA tasks. These results may be caused by the semantic similarity from CLIP models that are affected by complex sentences.
4.4.3 Comparison with Other Methods
Comparison with BiRD + RoPE-Mixed [12]. The RoPE-Mixed uses frequencies for both axes as learnable network parameters, effectively handling image features in diagonal direction. It achieves the best performance in different visual scenarios [48, 49, 23]. Hence we compare our MCG-RoPE with RoPE-Mixed. As shown in Table 3, by adding RoPE-Mixed to BiRD [14], great achievements take place on all modalities (CT, MR, X-ray, etc.). Compared with RoPE-Mixed [12], our MCG-RoPE outperforms it by 4.95%, 2.94%, 5.78% and 2.60% on VG, ROC, RC and MIA tasks, respectively. For more detailed performance on the 8 modalities, MCG-RoPE still achieves the best improvements most of the time.
Comparison with BiRD + $V_{k}D$ [30]. $V_{k}D$ is a novel method that can improve knowledge distillation. This work utilizes a projection layer for maximizing the knowledge transfer to the student backbone. Many experiments [8, 33, 31, 13] have been made to prove the best performance of this method. Compared with BiRD with $V_{k}D$ in the Table 3, Pseudo-KD shows its advantages on multi-task fine-grained Med-VQA scenarios, outperforming $V_{k}D$ almost on all modalities.
Figure 5: Visualization of attentions on GC and VG tasks. The ClinKD generates better answers compared with BiRD [14].
4.4.4 Case Study
Case study on Med-GRIT-Test30 [39, 7]. Figure 5 illustrates the visualization of the attention map and a comparison among BiRD [14], ClinKD (ours), and the ground truth. The grounding of ClinKD is more accurate because MCG-RoPE can improve image-text alignment, and at the same time, Pseudo-KD enables ClinKD to better adapt to medical knowledge.
Case study on Additional VQA-RAD [16]. As shown in Figure 6, we conducted an experiment on VQA-RAD [16]. Despite the relatively simple annotations, ClinKD provides fine-grained answers with three accurate groundings, demonstrating its precision in response.
To assess the models’ attention to pathological features, we conducted a comparative analysis of three RoPE variants. The results indicate that 2D-RoPE [32] tends to show low attention on pathological regions. In contrast, RoPE-Mixed [12] demonstrates a greater emphasis on diagonal features. Notably, our Med-CLIP Guided RoPE specifically targets and enhances the model’s attention to pathological regions, thereby facilitating more accurate alignment between text and images, especially in the identification of relevant medical features.
Figure 6: An example from VQA-RAD [16] and the performance of RoPE variants.
4.4.5 Limitation
Although our method alleviates, to some extent, the issue of insufficient prior knowledge and achieves impressive state-of-the-art performance on the Med-GRIT-270k [39, 7] dataset, the model still exhibits hallucinations. Furthermore, through comparative evaluations on the LLaVA-Med-qa0.2k [46] dataset, our approach remains insufficient to enable the model to emerge with strong generalization capabilities in Med-VQA domain. Creating high-quality Med-VQA datasets and training medical prior knowledge before freezing the visual encoder could significantly enhance the model’s generalization ability on medical data.
5 Conclusion
In this work, we propose ClinKD, a novel medical distiller centered on Pseudo-Labels Medical Knowledge Distillation and Reflective Correction Training. In addition, we propose the Med-CLIP Guided RoPE to improve image-text alignment. They are designed to enhance the performance on fine-grained multi-task datasets in Med-VQA domain. Extensive experiments demonstrate that our Med-CLIP Guided RoPE achieves superior image-text alignment, while the Pseudo-Labels Medical Knowledge Distillation effectively bridges prior medical knowledge gaps. These mechanisms synergistically enhance the model’s medical knowledge adaptation capabilities.
In future work, further refinements to the distillation process, along with the incorporation of more comprehensive medical knowledge and higher-quality datasets, could significantly improve model’s robustness and reduce the model’s reliance on manually labeled data.
6 Acknowledgments
This work was supported by the Natural Science Foundation of Guangdong Province (No. 2023A1515010673), in part by the Shenzhen Science and Technology Innovation Bureau key project (No. JSGG20220831110400001, No. CJGJZD20230724093303007,KJZD20240903101259001), in part by Shenzhen Medical Research Fund (No. D2404001), in part by Shenzhen Engineering Labora- tory for Diagnosis & Treatment Key Technologies of Interventional Surgical Robots (XMHT20220104009), and the Key Laboratory of Biomedical Imaging Science and System, CAS, for the Research platform support.
Supplementary Material
A1 Datasets in the Experiments
A1.1 Med-GRIT-270k [14]
Instance-level Meta Information. Initially, approximately 60,000 biomedical image-mask pairs were randomly sampled from the SA-Med2D-20M dataset  [39, 7], considering modality diversity and redundancy. Notably, the original dataset contains numerous 2D slices extracted from 3D volumes, which could lead to high similarity and redundancy among images. Next, instance locations were determined based on instance-level masks. Specifically, spatial coordinates of each instance were represented in textual form as $[X_{topleft},Y_{topleft},X_{bottomright},Y_{bottomright}]$, and subsequently normalized to the range [0, 1]. Finally, each image was annotated with additional meta-information, including imaging modality, scanned anatomical region, orientation, and instance coordinates.
Obtaining Image Captions. Carefully designed prompts, combined with the meta-information, were provided to ChatGPT to extract global context information for each image.
Details of Instruction-Tuning Data. Spatial understanding is manifested through various task formats. This primarily encompasses two distinct types and their corresponding task names: (i) Region-in and Text-out: Referring Object Classification (ROC), Referring Captioning (RC), (ii) Text-in and Region-out: Visual Grounding (VG), and (iii) Text-in and Text-out: Medical Image Analysis (MIA). To reduce ambiguity and enhance the model’s capability for fine-grained visual comprehension, some essential strategies are adopted. The special tokens (<ref >and </ref>) are introduced, marking the content referred to by the bounding box. This aptly associates bounding boxes with their corresponding descriptive words or sentences. Subsequently, we instructed ChatGPT to design a question and answer for each task. Finally, We mapped the coordinates within the range [0, 1000] and reformatted them as $(X_{topleft},Y_{topleft})$, $(X_{bottomright},Y_{bottomright})$. To differentiate between detection strings and regular text strings, two special tokens (<box>and </box>) are appended at the start and end of the bounding box string, respectively.
The proportions of various modalities are shown in the Figure A1 and the Figure A2 shows an example of instruction-tuning conversations.
Figure A1: The proportions of various modalities in the Med-GRIT-270k [14]
A1.2 LLaVA-Med-qa0.2k [18]
The LLaVA-Med-qa0.2k is primarily used for evaluating LLaVA-Med [18]. It is an evaluation dataset consisting of 193 novel questions was constructed. Specifically, 50 previously unseen image-caption pairs were randomly selected from the PMC-15M dataset. Conversational questions were produced using the same self-instruct data generation pipeline employed in the second training stage. In contrast, detailed description questions were randomly chosen from a predefined question set, aiming to prompt comprehensive and detailed responses.
Figure A2: The example of instruction-tuning data.
A2 Training Configurations
We use LLaMA-Factory to train Qwen2-VL and here are some information about model configurations and training configurations.
Category Key Value
Formatting ShareGPT
Columns messages messages
images images
Tags role_tag role
content_tag content
user_tag user
assistant_tag assistant
Table A1: Qwen2-VL [34] Data Configuration Summary.
Parameter Value
Model Architecture
Hidden size 3584
Hidden layers 28
Attention heads 28
Key/Value heads 4
Intermediate size 18944
Max position embeddings 32768
Vocab size 152,064
RMS norm epsilon 1e-06
Training/Regularization
Initializer range 0.02
Attention dropout 0.0
Sliding window size 32768
Special Tokens
BOS token ID 151643
EOS token ID 151645
Image token ID 151655
Video token ID 151656
Vision start token ID 151652
Vision end token ID 151653
Vision Configuration
Input channels 3
Spatial patch size 14
Table A2: The Architecture and Configuration Parameters of ClinKD.
Parameter Value
Model Configuration
Base model Qwen2-VL-7B-Instruct
Quantization method bitsandbytes
Fine-tuning type Full
Template qwen2_vl
Training Setup
Batch size 2
Gradient accumulation 8
Max sequence length 2048
Training epochs 3
Learning rate $2\times 10^{-5}$
Warmup steps 0
Optimizer AdamW
Learning schedule Cosine
Memory Management
DeepSpeed stage 3
Offloading Disabled
BF16 computation Enabled
Max grad norm 1.0
Adapter Configuration
LoRA rank 8
LoRA alpha 16
LoRA dropout 0
Trainable layers 2
Advanced Options
BAdam mode Layer (ascending)
BAdam update ratio 0.05
GaLoRe rank 16
SFT stage Supervised Fine-Tuning
Reward modeling Disabled
Table A3: Training Configuration Parameters for ClinKD.
During the knowledge distillation process, additional GPU memory allocation is required to accommodate the teacher model in inference mode. For instance, the Qwen2-VL-7B model consumes approximately 25GB of GPU memory per question-answering iteration. When processing large batches of data, we recommend allocating at least 80GB of GPU memory for the teacher model to ensure smooth operation.