### v4

# ClinKD: Cross-Modal Clinical Knowledge Distiller For Multi-Task Medical Images

Abstract
Medical Visual Question Answering (Med-VQA) represents a critical and challenging subtask within the general VQA domain. Despite significant advancements in general VQA, multimodal large language models (MLLMs) still exhibit substantial limitations when handling multi-task VQA scenarios. These limitations manifest through erroneous spatial localization and misinterpretation of medical images, which primarily arise from two fundamental issues: inadequate image-text alignment and insufficient domain-specified knowledge for medical applications. To address these issues, we introduce the Cross-Modal Clinical Knowledge Distiller (ClinKD), an innovative framework designed to enhance image-text alignment and establish more effective medical knowledge transformation mechanisms, which enables MLLMs to perform better even when lacking prior medical knowledge. Our extensive experimental evaluations demonstrate that the ClinKD achieves state-of-the-art performance on several datasets which are challenging for the Med-VQA task. The results indicate that our approach not only significantly improves image-text alignment but also effectively enables MLLMs to adapt to the medical knowledge. The source code for ClinKD is available at: https://github.com/overloadedHenry/ClinKD. Refer to caption Figure 1: The left part shows that the BiRD huang2024BiRD only utilizes traditional supervised fine-tuning (SFT) and 2D-RoPE su2023roformerenhancedtransformerrotary to obtain the capabilities of understanding and grounding medical images. The right part is our method which uses Med-CLIP Guided RoPE for enhancing image-text alignment and Adaptive Confidence-Margin Curriculum Pseudo-KD for filling the gap of medical knowledge in ClinKD. During the SFT procedure, we will use Reflective Correction Training to enable answers with low semantic similarity or incorrect groundings to be reused.
1 Introduction
Medical Visual Question Answering (Med-VQA) Lin_2023 ; Abacha2019VQAMedOO ; Hasan2018OverviewOI is a critical domain for applying MLLMs yang2024modelmergingllmsmllms ; liang2024comprehensivesurveyguidemultimodal ; caffagni-etal-2024-revolution to medical image analysis li2024towards . Recent advancements in MLLMs have enabled them to achieve preliminary capabilities in image analysis, annotation, and user instruction compliance chen2025rllavaimprovingmedvqaunderstanding . Previous studies have made multiple attempts yang2025llmmedqaenhancingmedicalquestion ; kumar2024medvisionllamaleveragingpretrainedlarge ; he2024parameterefficientfinetuningmedicalmultimodal ; liu2024hcllmhistoricalconstrainedlargelanguage ; li2024llava ; huang2024BiRD to enhance the performance of general-purpose MLLMs in Med-VQA.
For instance, LLaVA-Med li2024llava extends instruction-tuning techniques for MLLMs to the medical domain. This work applied instruction tuning and curated image-text datasets to align visual-textual features in LLaVA llava , enabling LLaVA-Med to recognize and analyze medical images. However, this system lacked fine-grained image understanding, causing erroneous spatial localization when questioned about specific pathological locations gai2024enhancing ; huang2024BiRD .
Subsequently, the BiRD huang2024BiRD , a model that incorporated fine-grained image recognition by developing a biomedical refer-and-ground instruction-tuning dataset, addresses the absence of grounding capabilities you2023ferretrefergroundgranularity ; zhang2024ferretv2improvedbaselinereferring in LLaVA-Med’s work li2024llava . Nevertheless, as shown in Figure 1, due to its reliance on traditional supervised fine-tuning (instruction-tuning) sft1 ; sft2 ; sft3 and conventional 2D Rotary Position Embedding (RoPE) su2023roformerenhancedtransformerrotary , BiRD tends to generate erroneous spatial localization and misinterpretation of medical images because of incomplete image-text alignment alignment1 ; alignment2 . Moreover, the visual encoder of BiRD is constrained by the lack of medical knowledge, resulting in difficulties in medical knowledge adaptation adaptation ; lin2025healthgptmedicallargevisionlanguage .
To overcome these limitations, we propose a medical distillation framework named ClinKD to boost capabilities in the Med-VQA domain. Specifically, our method mainly focuses on two issues: incomplete image-text alignment and the professional medical knowledge gap between the general-purpose MLLMs and medical domain that requires finer-grained medical knowledge medical_fine-grained ; zeng2024visualorientedfinegrainedknowledgeediting .
For the first issue, we propose Med-CLIP Guided Rotary Position Embedding (MCG-RoPE), a novel position embedding method that focuses on the inter-modal and intra-modal intervals during joint image-text training, while ensuring the equivalence of context intervals immediately adjacent to images. Compared to traditional RoPE su2023roformerenhancedtransformerrotary , the MCG-RoPE uses the distinct index intervals to better capture cross-modal information, thereby achieving more effective image-text alignment.
For the second issue, we propose Clinical Knowledge Distiller (ClinKD) which consists of Adaptive Confidence-Margin Curriculum Pseudo-KD (Pseudo-KD) and Reflective Correction Training. Pseudo-KD provides prior medical knowledge by using pseudo-labels pseudolabels ; chen2023mixedpseudolabelssemisupervised , filling the gap of prior medical knowledge so that the model can better adapt to medical knowledge during the supervised fine-tuning. Reflective Correction Training allows samples with low semantic similarity to be sent for training again after being enhanced by MLLMs.
Our contributions are summarized as follows:
  • We introduce the ClinKD, an innovative distiller designed to establish more effective medical knowledge adaptation mechanisms. The code and checkpoint will be released.
  • For solving incomplete image-text alignment, we propose the Med-CLIP Guided Rotary Position Embedding (MCG-RoPE). The MCG-RoPE utilizes the distinct inter-modal and intra-modal features to improve image-text alignment.
  • For bridging medical knowledge gap between general-purpose MLLMs and specialized medical applications, we propose the Adaptive Confidence-Margin Curriculum Pseudo-KD, filling the medical knowledge gap before supervised fine-tuning.
  • Extensive experimental results show that the proposed ClinKD achieves competitive or superior performance on test set of diverse tasks and datasets.
Figure 2: Overview of our ClinKD’s framework: The framework begins with multi-task medical images being segmented into patches. The med-extractor module with Med-CLIP Guided Rotary Position Embedding extracts dimensional features from the patched image embeddings. The ClinKD system undergoes distillation before supervised fine-tuning (SFT). During the SFT, answers with low semantic similarity or incorrect groundings are prioritized for retraining. When responding to queries, the framework selects the answer with the highest score, ensuring accurate medical image analysis and interpretation.
2 Related Work
2.1 Biomedical Instruction-Tuning Datasets
Many Med-VQA datasets such as VQA-RAD lau2018dataset and SLAKE liu2021slake contain only medical images and simple QA pairs. Specifically, these QA pairs do not have the fine-grained descriptions of pathological features. This limitation constrains the ability of MLLMs to perform accurate spatial localization. To boost and evaluate MLLMs’ capabilities in the Med-VQA domain, many biomedical instruction-tuning datasets are designed.
For example, the LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation contains 50 medical images sampled from several modalities such as CT, Angiography and PET and 193 questions according to these medical images. The medical QA pairs are generated by GPT-4 based on the metadata. This dataset is designed to evaluate the capabilities of describing medical images, specifically the biomedical features.
Unlike the LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation , the Med-GRIT-270k huang2024BiRD includes multi-task fine-grained medical QA pairs especially for grounding spatial locations. It is designed to enhance the model’s capabilities in referring and grounding. This multi-task dataset comprises four distinct tasks: Visual Grounding (VG), Referring Object Classification (ROC), Referring Captioning (RC), and Medical Image Analysis (MIA). The VG task evaluates the model’s capability of accurately matching text descriptions to corresponding image regions. The RC task assesses the model’s capability to recognize specific image areas and generate descriptive captions for them. The ROC task examines the model’s understanding of textual information related to particular image regions and their associated visual details. The MIA task evaluates the model’s comprehension of medical images and their multi-modal context. All data is sourced from eight modalities: CT, MR, X-ray, PET, Endoscopy, Dermoscopy, Fundus, and Ultrasound. Utilizing data from these eight modalities allows for a more comprehensive evaluation of the model’s capabilities.
2.2 Multimodal Large Language Models (MLLMs) for Biomedicine
With the development of MLLMs, an increasing number of approaches have been applied to the medical field, such as BioMedGPT Zhang_2024 , LLaVA-Med li2024llava , and BiRD huang2024BiRD . These methods drive the development of MLLMs in the biomedical domain from different perspectives. For instance, LLaVA-Med li2024llava applied instruction fine-tuning to the medical imaging field and built a medical question-answering system based on LLaVA llava , which provided reasonable responses by processing user-submitted images and instructions. However, LLaVA-Med did not support fine-grained interactions with medical images, meaning that the system cannot precisely locate pathological regions within the images. To address the lack of referring and grounding capabilities, the BiRD huang2024BiRD , which is based on Qwen-VL bai2023qwenvlversatilevisionlanguagemodel , leveraged a finer-grained medical dataset. However, the ViT of BiRD was not trained by sufficient medical data before being frozen, resulting in a gap in medical knowledge domain. Compared with their work huang2024BiRD ; li2024llava , our method modifies the position embedding method and introduces an effective Confidence-Margin Curriculum Pseudo-KD framework.
2.3 Rotary Position Embedding (RoPE)
The introduced RoPE su2023roformerenhancedtransformerrotary makes the relative position features of the context in the text more easily captured by LLMs. Further, the RoPE-Mixed heo2024rotarypositionembeddingvision introduced mixed learnable frequencies to RoPE, enabling it to handle diagonal directions and making RoPE itself learnable. It showed more efficiency when dealing with diagonal features of the image. The VideoRoPE wei2025videoropemakesgoodvideo extended the diagonal layout and variable frequency to the video domain, demonstrating its advantages in video analysis tasks. Unlike existing methods, our approach focuses on 2D multi-task medical images where capturing semantic information is challenging. We utilize cross-modal intervals to enable the model to better distinguish modal information, thereby improving image-text alignment.
3 Methodology
Figure 3: The visualized example of Med-CLIP Guided Rotary Position Embedding. The original vector will be rotated by different angles according to the transformed position indexes. The inter-modal and intra-modal intervals are different so that the Med-CLIP Guided Rotary Position Embedding can better align image feature and text feature.
As shown in Figure 2, we aim at enhancing the ClinKD’s performance by mainly leveraging Med-CLIP Guided Rotary Position Embedding, Pseudo-KD, Reflective Correction Training and Semantic-Aware Selective Generation. The proposed Med-CLIP Guided Rotary Position uses the distinct inter-modal and intra-modal space to improve image-text alignment. The distiller provides prior medical knowledge to enhance the model’s capability of medical knowledge adaptation. In the inference procedure, the Semantic-Aware Selective Generation will allow the model to choose the answer with the best score.
3.1 Med-CLIP Guided Rotary Position Embedding
The standard 2D Rotary Position Embedding (RoPE) su2023roformerenhancedtransformerrotary does not explicitly account for cross-modal relationships between image and text tokens. To address this, we introduce the Med-CLIP Guided Rotary Position Embedding (MCG-RoPE), which deal with tokens from different modals by considering splited queries to improve image-text alignment. Figure 3 illustrates an example of our approach.
Distinct Inter-Modal Intervals.
Consider a sequence of position indices for a text–image–text pattern:
$S_{\mathrm{idx}}=\bigl{(}T_{1},T_{2},\dots,T_{\ell},\;V_{\ell+1},\dots,V_{\ell+wh},\;T_{\ell+wh+1},\dots,T_{\ell+wh+k}\bigr{)},$ (1)
where each text index $T_{i}=(t_{i},t_{i})\in\mathbb{R}^{2}$ and each image index $V_{j}=(v^{(1)}_{j},v^{(2)}_{j})\in\mathbb{R}^{2}$. We apply an affine transform $\mathcal{F}$ to each image index so that intra-modal and inter-modal gaps become equal along both axes:
$\mathcal{F}\bigl{(}v_{i}^{(1)},v_{i}^{(2)}\bigr{)}=\bigl{(}\alpha_{1}\,v_{i}^{(1)}+\lambda_{1},\;\alpha_{2}\,v_{i}^{(2)}+\lambda_{2}\bigr{)}=:\mathcal{V}_{i},$ (2)
with scaling factors $\alpha_{1},\alpha_{2}>0$ and offsets $\lambda_{1},\lambda_{2}\in\mathbb{R}$. By enforcing a pronounced interval difference between text and image tokens, the model is guided to focus on fine-grained visual details in medical images while maintaining clear separation from surrounding text.
Distinct Interval-Based Rotation.
Let $q\in\mathbb{R}^{d_{h}+d_{w}}$ be a query/key vector, split as
$q=\bigl{[}q^{(h)};q^{(w)}\bigr{]},\quad q^{(h)}\in\mathbb{R}^{d_{h}},\;q^{(w)}\in\mathbb{R}^{d_{w}}.$ (3)
By processing the height and width dimensions separately, the model can capture both coarse inter-modal alignment and fine intra-modal structure.
Define frequency scales for each half of the embedding:
$\omega_{i}^{(h)}=10\,000^{-\frac{2i}{d_{h}}},\quad\omega_{i}^{(w)}=10\,000^{-\frac{2i}{d_{w}}},\quad i=1,\dots,\tfrac{d_{h}}{2}\;\text{or}\;\tfrac{d_{w}}{2}.$ (4)
Given the transformed positions $x_{h}$ and $x_{w}$, the rotation angles satisfy
$\theta_{x_{h},i}^{(h)}=x_{h}\,\omega_{i}^{(h)},\quad\theta_{x_{w},i}^{(w)}=x_{w}\,\omega_{i}^{(w)}.$ (5)
We construct block-diagonal rotation matrices
$\mathbf{R}_{\Theta,x_{h}}^{d_{h}}=\mathrm{diag}\bigl{(}R_{\theta_{x_{h},1}^{(h)}},\dots,R_{\theta_{x_{h},\frac{d_{h}}{2}}^{(h)}}\bigr{)},\quad\mathbf{R}_{\Theta,x_{w}}^{d_{w}}=\mathrm{diag}\bigl{(}R_{\theta_{x_{w},1}^{(w)}},\dots,R_{\theta_{x_{w},\frac{d_{w}}{2}}^{(w)}}\bigr{)},$ (6)
where each $2\times 2$ block is
$R_{\theta}=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}.$ (7)
Finally, the transformed vector is
$\widetilde{q}=\bigl{[}\mathbf{R}_{\Theta,x_{h}}^{d_{h}}\,q^{(h)};\,\mathbf{R}_{\Theta,x_{w}}^{d_{w}}\,q^{(w)}\bigr{]}.$ (8)
By emphasizing distinct interval gaps and splitting the query into separate height and width components, MCG-RoPE enables the model to more effectively capture fine-grained features in medical images and achieve precise image–text alignment. The difference between 2D-RoPE and MCG-RoPE is shown in Appendix.
3.2 Adaptive Confidence-Margin Curriculum Pseudo-KD
In order to mitigate the negative impact of noisy pseudo-labels and more effectively guide the student model, we propose an Adaptive Confidence–Margin Curriculum Pseudo-KD strategy. This approach extends standard pseudo-label distillation by (1) weighting each sample’s distillation loss according to both the teacher’s confidence and the margin between top logits, and (2) employing a simple curriculum learning schedule that gradually relaxes the confidence threshold over training.
Confidence and Margin Computation.
Given an input example $x$, the teacher network produces logits $z_{t}=f_{t}(x)$. We compute the softmax probabilities
$p_{t_{i}}(x)\;=\;\frac{\exp\bigl{(}z_{t_{i}}\bigr{)}}{\sum_{j}\exp\bigl{(}z_{t_{j}}\bigr{)}}\,.$ (9)
We define the confidence and margin measures as
$C(x)\;=\;\max_{i}p_{t_{i}}(x)\quad,\quad\Delta(x)\;=\;p_{t_{(1)}}(x)\;-\;p_{t_{(2)}}(x)\,,$ (10)
where $p_{(1)}$ and $p_{(2)}$ denote the largest and second-largest probabilities, respectively.
Curriculum Threshold Schedule.
Let $t\in[0,1]$ denote the normalized training progress (i.e., fraction of total steps completed). We set an initial confidence threshold $\tau_{0}$ and a minimum threshold $\tau_{\min}$. The threshold decays linearly:
$\tau(t)\;=\;\tau_{\min}\;+\;(\tau_{0}-\tau_{\min})\,\bigl{(}1-t\bigr{)}\,.$ (11)
Adaptive Weighting Function.
We combine confidence and margin into a per-sample weight:
$w(x)\;=\;\begin{cases}\bigl{[}C(x)\bigr{]}^{\gamma}\,\bigl{[}\Delta(x)\bigr{]}^{\beta},&C(x)\;\geq\;\tau(t)\,,\\ 0,&\text{otherwise},\end{cases}$ (12)
where $\gamma,\beta\in(0,1]$ control sensitivity to confidence and margin.
Weighted Distillation Loss.
With student logits $z_{s}$ and temperature $T$, the weighted KL-divergence distillation loss is
$\mathcal{L}_{\mathrm{KD}}=\frac{1}{N}\sum_{i=1}^{N}w(x_{i})\;\mathbb{D}_{KL}\Bigl{(}softmax\bigl{(}z_{t_{i}}/T\bigr{)}\,\big{\|}\,softmax\bigl{(}z_{s_{i}}/T\bigr{)}\Bigr{)}\,.$ (13)
Overall Objective.
We combine the standard cross-entropy loss $L_{\mathrm{CE}}$ with our adaptive pseudo-KD term:
$\mathcal{L}_{Distill}\;=\;\dfrac{\mathcal{L}_{\mathrm{CE}}\;+\;\alpha\,\mathcal{L}_{\mathrm{KD}}}{1+\alpha}\quad,\quad\alpha\in[0,1]\,.$ (14)
3.3 Reflective Correction Training
In this part, during each round of supervised fine-tuning, we record the model’s prediction $\hat{\mathbf{y}}=(\mathbf{t}_{o},\mathbf{c})$ and the corresponding ground-truth label $\mathbf{y}=(\mathbf{t}_{gt},\mathbf{c})$. We then compute a weighted cosine similarity to assess semantic alignment and, if necessary, invoke GPT-4o to “correct” the output in order to enrich label diversity. The procedure is as follows:
First, map the text labels into a $d$-dimensional embedding space to obtain $\mathbf{e}_{o},\mathbf{e}_{gt}\in\mathbb{R}^{d}$, while retaining the shared anchor-box coordinate vector $\mathbf{c}\in\mathbb{R}^{4}$.
Next, define the weighted cosine similarity as
$S(\mathbf{e}_{o},\mathbf{e}_{gt};\mathbf{c})=\frac{\mu\,\langle\mathbf{e}_{o},\mathbf{e}_{gt}\rangle+\nu\,\langle\mathbf{c},\mathbf{c}\rangle}{\sqrt{\bigl{(}\mu\,\|\mathbf{e}_{o}\|^{2}+\nu\,\|\mathbf{c}\|^{2}\bigr{)}\,\bigl{(}\mu\,\|\mathbf{e}_{gt}\|^{2}+\nu\,\|\mathbf{c}\|^{2}\bigr{)}}}\,,$ (15)
[…]
We then compare $S$ against $\tau$ and decide whether to trigger GPT-4o correction:
$\hat{\mathbf{y}}^{\prime}=\begin{cases}\mathrm{GPT4o\_Correct}\bigl{(}\hat{\mathbf{y}},\,\mathbf{y}\bigr{)},&S(\mathbf{e}_{o},\mathbf{e}_{gt};\mathbf{c})<\tau,\\[6.0pt] \hat{\mathbf{y}},&S(\mathbf{e}_{o},\mathbf{e}_{gt};\mathbf{c})\geq\tau,\end{cases}$ (16)
[…]
3.4 Semantic-Aware Selective Generation
The responses generated by large language models exhibit randomness, and relying on a single inference may result in missing the optimal answer. Semantic-Aware Selective Generation means that instead of directly using the model’s single-round generation for evaluation, we allow the model to generate multiple samples. We then utilize the CLIP models to score the alignment between the generated text and image. Finally, we reorder the outputs based on these scores and select the one with the highest score as the final result. The pseudo code is shown in Algorithm.
4 Experiments
4.1 Datasets
We train ClinKD on the Med-GRIT-270k huang2024BiRD dataset, and evaluate the performance on Med-GRIT-30k and LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation dataset.
We alse test the few-shot performance on VQA-RAD lau2018dataset , SLAKE liu2021slake and Path-VQA he2020pathvqa . For more information, please refer to Appendix.
4.2 Evaluation Metrics
We use Recall@0.5 for Visual Grounding (VG), Recall for Referring Object Classification (ROC), SPICE anderson2016spicesemanticpropositionalimage for Referring Caption (RC), and mBMR huang2024BiRD for Medical Image Analysis (MIA). The final overall result is computed as the mean of the evaluation scores across these four tasks.
For the evaluation on LLaVA-Med-qa0.2k li2024llava which has 0.2k QA pairs of MIA task, we use mBMR as the metric.
In few-shot experiments, we use BLEU-1 and F1-score to evaluate models’ performance.
4.3 Analysis
4.3.1 Comparison with the State-of-the-Art Methods
Evaluation on the Med-GRIT-Test30k ye2023samed2d20m ; cheng2023sammed2d . As shown in Table 1, we compare ClinKD with BiRD huang2024BiRD and LLaVA-Med li2024llava . Results show that ClinKD achieves scores of 67.51%, 82.35%, 70.56%, and 65.69% in the VG, ROC, RC, and MIA tasks respectively, outperforming other comparisons by a clear margin in all tasks. On average, ClinKD outperforms BiRD huang2024BiRD and LLaVA-Med li2024llava by 14.87% and 66.00% respectively. ClinKD’s advantages lie in its more adequate image-text alignment and the prior medical knowledge provided by Pseudo-KD.
Evaluation on the LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation . Table 1 represents the models’ performance on LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation . Our ClinKD model demonstrates superior performance on this dataset, outperforming both LLaVA-Med (by 3.50%) and BiRD (by 2.48%). The relatively low mBMR scores can be attributed to the notable language style discrepancies observed between the training and testing datasets.
Model Test dataset VG (Recall@0.5) ROC (Recall) RC (SPICE) MIA (mBMR) Average
LLaVA-Med li2024llava Med-GRIT-Test30k ye2023samed2d20m ; cheng2023sammed2d 0 2.75 8.18 11.20 5.53
BiRD huang2024BiRD Med-GRIT-Test30k ye2023samed2d20m ; cheng2023sammed2d 53.92 65.33 55.23 52.17 56.66
ClinKD (Ours) Med-GRIT-Test30k ye2023samed2d20m ; cheng2023sammed2d 67.51 82.35 70.56 65.69 71.53
LLaVA-Med li2024llava LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation - - - 20.04 -
BiRD huang2024BiRD LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation - - - 21.06 -
ClinKD (Ours) LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation - - - 23.54 -
Table 1: Comparison with LLaVA-Med li2024llava and BiRD (previous SOTA) huang2024BiRD . The best scores are shown in bold.
Model Proposed Methods VG (Recall@0.5) ROC (Recall) RC (SPICE) MIA (mBMR) Average
MCG-RoPE Pseudo-KD SASG
Qwen2-VL wang2024qwen2vlenhancingvisionlanguagemodels 55.43 (+1.51) 68.33 (+3.00) 57.39 (+2.16) 62.23 (+10.06) 61.37 (+4.71)
BiRD huang2024BiRD 53.92 65.33 55.23 52.17 56.66
64.03 (+10.11) 72.74 (+7.41) 61.32 (+6.09) 59.63 (+7.46) 64.44 (+7.78)
66.33 (+12.41) 79.52 (+14.19) 67.21 (+11.98) 64.32 (+12.15) 69.35 (+12.69)
55.89 (+1.97) 67.84 (+2.51) 54.32 (-0.91) 51.27 (-0.90) 57.33 (+0.67)
ClinKD (Ours) 67.51 (+13.59) 82.35 (+17.02) 70.56 (+15.33) 65.69 (+13.52) 71.53 (+14.87)
Table 2: Ablation study for different models and proposed methods. The value which is blue shows how much the metric scores improved. The best scores are shown in bold. The MCG-RoPE, Pseudo-KD and SASG stand for Med-CLIP Guided RoPE, Adaptive Confidence-Margin Curriculum Pseudo-KD and Semantic-Aware Selective Generation respectively.
4.3.2 Ablation study
Effect of Qwen2-VL wang2024qwen2vlenhancingvisionlanguagemodels . Table 2 presents the ablation study of proposed methods. To investigate the impact of model upgrade on the experiment, we applied the same training strategy to Qwen2-VL as we did to BiRD huang2024BiRD . The results show that the model upgrade only has a significant effect on the MIA task, which shows Qwen2-VL wang2024qwen2vlenhancingvisionlanguagemodels outperforms BiRD by 10.06%.
Effect of Med-CLIP Guided Rotary Position Embedding. As shown in Table 2, with adding MCG-RoPE to BiRD, all metrics exhibit significant enhancements, with the average rising from 56.66% to 64.44%. This indicates the importance of adequate image-text alignment.
Effect of Adaptive Confidence-Margin Curriculum Pseudo-KD. Table 2 represents improvements achieved by using Pseudo-KD that leads to enhancements by 12.41%, 14.19%, 11.98% and 12.15% in the VG, ROC, RC and MIA tasks, respectively. The results illustrate that providing prior medical knowledge by using Pseudo-KD can fill MLLMs’ gap of medical knowledge, boosting capabilities of adapting to new medical knowledge. Figure shows the effect of different values of $\alpha$ in Eq. (14). In this experiment, the average performance achieves the best when we set $\alpha=0.5$. Both excessively high and low alpha values can lead to performance degradation. This might be because a low $\alpha$ value prevents ClinKD from learning sufficient prior medical knowledge, while an overly high alpha value tends to make ClinKD generate irrational pseudo-labels that contradict common medical knowledge.
Effect of Semantic-Aware Selective Generation. From the results in Table 2, we can see a light enhancement on VG and ROC task. However, there is a slight decrease in the metric scores on the RC and MIA tasks. These results may be caused by the semantic similarity from CLIP models that are affected by complex sentences.
Task Metric Method CT MR X-ray PET Endoscopy Dermoscopy Fundus Ultrasound Average
BiRD huang2024BiRD 44.47$\pm$0.14 29.26$\pm$0.11 41.73$\pm$0.16 56.46$\pm$0.15 53.60$\pm$0.12 75.63$\pm$0.17 84.15$\pm$0.13 46.04$\pm$0.18 53.92$\pm$0.14
BiRD huang2024BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision 46.90$\pm$0.16 53.11$\pm$0.13 42.03$\pm$0.17 60.84$\pm$0.14 53.49$\pm$0.15 78.25$\pm$0.11 85.16$\pm$0.12 52.83$\pm$0.13 59.08$\pm$0.19
BiRD huang2024BiRD + MCG-RoPE (Ours) 51.22$\pm$0.14 49.40$\pm$0.18 45.92$\pm$0.17 68.60$\pm$0.13 60.37$\pm$0.19 88.28$\pm$0.12 86.59$\pm$0.15 61.88$\pm$0.16 64.03$\pm$0.15
BiRD huang2024BiRD + $V_{k}D$  Miles_2024_CVPR 58.77$\pm$0.12 50.95$\pm$0.16 45.69$\pm$0.11 65.89$\pm$0.18 57.63$\pm$0.14 93.32$\pm$0.15 79.27$\pm$0.12 59.90$\pm$0.15 63.93$\pm$0.19
BiRD huang2024BiRD + Pseudo-KD (Ours) 62.59$\pm$0.15 52.86$\pm$0.17 47.19$\pm$0.13 66.28$\pm$0.16 59.84$\pm$0.11 93.68$\pm$0.18 85.37$\pm$0.14 62.87$\pm$0.19 66.33$\pm$0.13
VG Recall@0.5 ClinKD (Ours) 65.52$\pm$0.13 52.23$\pm$0.30 48.56$\pm$0.19 69.25$\pm$0.12 60.37$\pm$0.16 94.22$\pm$0.10 86.59$\pm$0.14 64.47$\pm$0.15 67.51$\pm$0.18
BiRD huang2024BiRD 34.76$\pm$0.11 61.79$\pm$0.15 53.74$\pm$0.14 - 60.40$\pm$0.17 96.61$\pm$0.18 - 84.65$\pm$0.12 65.33$\pm$0.14
BiRD huang2024BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision 39.78$\pm$0.13 62.13$\pm$0.19 74.27$\pm$0.11 - 61.33$\pm$0.13 96.61$\pm$0.17 - 84.65$\pm$0.15 69.80$\pm$0.12
BiRD huang2024BiRD + MCG-RoPE (Ours) 44.47$\pm$0.15 64.20$\pm$0.18 83.75$\pm$0.11 - 62.19$\pm$0.14 96.61$\pm$0.19 - 85.22$\pm$0.16 72.74$\pm$0.13
BiRD huang2024BiRD + $V_{k}D$ Miles_2024_CVPR 53.19$\pm$0.17 69.10$\pm$0.14 90.84$\pm$0.19 - 61.71$\pm$0.10 100$\pm$0.00 - 81.15$\pm$0.12 76.00$\pm$0.15
BiRD huang2024BiRD + Pseudo-KD (Ours) 58.47$\pm$0.16 69.92$\pm$0.12 95.83$\pm$0.18 - 66.32$\pm$0.11 100$\pm$0.00 - 86.58$\pm$0.13 79.52$\pm$0.14
ROC Recall ClinKD (Ours) 61.42$\pm$0.17 71.30$\pm$0.19 97.00$\pm$0.18 - 68.84$\pm$0.12 100$\pm$0.00 - 95.54$\pm$0.11 82.35$\pm$0.15
BiRD huang2024BiRD 41.88$\pm$0.12 51.69$\pm$0.14 37.39$\pm$0.18 47.95$\pm$0.15 54.07$\pm$0.10 77.44$\pm$0.16 48.73$\pm$0.13 82.65$\pm$0.19 55.23$\pm$0.14
BiRD huang2024BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision 43.51$\pm$0.15 51.88$\pm$0.18 54.61$\pm$0.13 52.83$\pm$0.19 54.07$\pm$0.11 68.15$\pm$0.17 53.99$\pm$0.12 65.28$\pm$0.14 55.54$\pm$0.13
BiRD huang2024BiRD + MCG-RoPE (Ours) 50.73$\pm$0.14 53.11$\pm$0.16 60.33$\pm$0.12 70.16$\pm$0.18 54.07$\pm$0.11 70.18$\pm$0.19 61.52$\pm$0.15 70.46$\pm$0.17 61.32$\pm$0.13
BiRD huang2024BiRD + $V_{k}D$  Miles_2024_CVPR 51.33$\pm$0.13 54.39$\pm$0.17 62.83$\pm$0.19 74.31$\pm$0.14 54.07$\pm$0.15 76.23$\pm$0.12 67.89$\pm$0.18 72.15$\pm$0.11 64.15$\pm$0.16
BiRD huang2024BiRD + Pseudo-KD (Ours) 52.78$\pm$0.12 58.41$\pm$0.15 70.39$\pm$0.18 75.85$\pm$0.13 55.22$\pm$0.16 81.34$\pm$0.11 71.54$\pm$0.17 72.15$\pm$0.14 67.21$\pm$0.19
RC SPICE anderson2016spicesemanticpropositionalimage ClinKD (Ours) 53.94$\pm$0.14 58.41$\pm$0.12 71.47$\pm$0.18 79.91$\pm$0.15 52.10$\pm$0.11 83.92$\pm$0.19 84.51$\pm$0.13 80.19$\pm$0.16 70.56$\pm$0.17
BiRD huang2024BiRD 47.01$\pm$0.14 49.35$\pm$0.18 37.17$\pm$0.15 57.15$\pm$0.12 39.91$\pm$0.17 72.13$\pm$0.13 48.87$\pm$0.16 65.78$\pm$0.19 52.17$\pm$0.11
BiRD huang2024BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision 57.10$\pm$0.17 50.89$\pm$0.13 42.55$\pm$0.16 69.28$\pm$0.11 41.63$\pm$0.18 73.32$\pm$0.14 59.63$\pm$0.15 61.80$\pm$0.10 57.03$\pm$0.19
BiRD huang2024BiRD + MCG-RoPE (Ours) 60.53$\pm$0.16 53.77$\pm$0.11 54.10$\pm$0.18 62.01$\pm$0.12 43.88$\pm$0.15 75.33$\pm$0.14 62.89$\pm$0.19 64.53$\pm$0.13 59.63$\pm$0.17
BiRD huang2024BiRD + $V_{k}D$  Miles_2024_CVPR 62.33$\pm$0.10 53.77$\pm$0.15 52.27$\pm$0.19 62.01$\pm$0.16 47.29$\pm$0.13 75.33$\pm$0.17 68.04$\pm$0.12 70.31$\pm$0.18 61.42$\pm$0.14
BiRD huang2024BiRD + Pseudo-KD (Ours) 66.85$\pm$0.11 55.73$\pm$0.14 52.27$\pm$0.17 64.74$\pm$0.15 50.58$\pm$0.12 78.21$\pm$0.19 70.02$\pm$0.13 76.16$\pm$0.18 64.32$\pm$0.16
MIA mBMR huang2024BiRD ClinKD (Ours) 66.85$\pm$0.13 56.82$\pm$0.16 54.01$\pm$0.18 67.08$\pm$0.14 50.58$\pm$0.11 81.00$\pm$0.15 70.02 $\pm$0.17 76.05$\pm$0.19 65.69$\pm$0.12
BiRD huang2024BiRD 43.03$\pm$0.12 48.02$\pm$0.11 42.51$\pm$0.17 53.85$\pm$0.19 51.99$\pm$0.14 80.45$\pm$0.16 60.58$\pm$0.13 69.78$\pm$0.15 -
BiRD huang2024BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision 46.82$\pm$0.17 54.50$\pm$0.12 53.37$\pm$0.15 60.98$\pm$0.13 52.63$\pm$0.18 79.08$\pm$0.10 66.26$\pm$0.19 66.14$\pm$0.11 -
BiRD huang2024BiRD + MCG-RoPE (Ours) 51.74$\pm$0.13 55.12$\pm$0.19 61.03$\pm$0.12 66.92$\pm$0.17 55.13$\pm$0.14 82.60$\pm$0.15 70.33$\pm$0.11 70.52$\pm$0.18 -
BiRD huang2024BiRD + $V_{k}D$ Miles_2024_CVPR 56.41$\pm$0.16 57.05$\pm$0.11 62.91$\pm$0.19 67.40$\pm$0.14 55.18$\pm$0.18 86.22$\pm$0.10 71.73$\pm$0.15 70.88$\pm$0.13 -
BiRD huang2024BiRD + Pseudo-KD (Ours) 60.17$\pm$0.14 59.23$\pm$0.16 66.42$\pm$0.12 68.96$\pm$0.19 57.99$\pm$0.13 88.31$\pm$0.15 75.64$\pm$0.17 74.44$\pm$0.10 -
Average - ClinKD (Ours) 61.93$\pm$0.15 59.69$\pm$0.18 67.76$\pm$0.19 72.08$\pm$0.13 57.97$\pm$0.16 89.79$\pm$0.11 80.37$\pm$0.14 79.06$\pm$0.12 -
Table 3: Comparison with other SOTA methods on Med-GRIT-Test30k ye2023samed2d20m ; cheng2023sammed2d . We evaluate the performance on eight modalities.
4.3.3 Comparison with Other Methods
Comparison with BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision . The RoPE-Mixed uses frequencies for both axes as learnable network parameters, effectively handling image features in diagonal direction. It achieves the best performance in different visual scenarios ade20k ; zhou2018semanticunderstandingscenesade20k ; coco . Hence we compare our MCG-RoPE with RoPE-Mixed. As shown in Table 3, by adding RoPE-Mixed to BiRD huang2024BiRD , great achievements take place on all modalities (CT, MR, X-ray, etc.). Compared with RoPE-Mixed heo2024rotarypositionembeddingvision , our MCG-RoPE outperforms it by 4.95%, 2.94%, 5.78% and 2.60% on VG, ROC, RC and MIA tasks, respectively. For more detailed performance on the 8 modalities, MCG-RoPE still achieves the best improvements most of the time.
Comparison with BiRD + $V_{k}D$ Miles_2024_CVPR . $V_{k}D$ is a novel method that can improve knowledge distillation. This work utilizes a projection layer for maximizing the knowledge transfer to the student backbone. Many experiments DLGAN_CVPR ; spkd ; pkd ; label-kd have been made to prove the best performance of this method. Compared with BiRD with $V_{k}D$ in the Table 3, Pseudo-KD shows its advantages on multi-task fine-grained Med-VQA scenarios, outperforming $V_{k}D$ almost on all modalities.
4.3.4 Few-shot Experiments
Table 4 shows the results of our additional experiments for testing few-shot performance. The evaluation scores shows that our ClinKD achieves the best performance on all datasets. It is can be seen that the BiRD which also utilizes Rotary Position Embedding outperforms LLaVA-Med li2024llava and Med-Flamingo med-fla . This may be caused by the special meaning of 2D-RoPE. For our ClinKD, the modified Med-CLIP Guided RoPE enhances the position information between image tokens and text tokens, enabling better image-text alignment.
Dataset Metric LLaVA-Med li2024llava BiRD huang2024BiRD Med-Flamingo med-fla ClinKD
VQA-RAD lau2018dataset F1 42.91 46.12 41.79 59.21
BLEU-1 43.35 48.23 43.27 57.09
SLAKE liu2021slake F1 44.28 49.25 45.86 58.64
BLEU-1 43.35 48.23 41.98 58.63
Path-VQA he2020pathvqa F1 44.28 49.25 42.56 59.21
Table 4: Comparison with LLaVA-Med li2024llava and BiRD (previous SOTA) huang2024BiRD . The best scores are shown in bold.
4.3.5 Case Study
Case study on Med-GRIT-Test30 ye2023samed2d20m ; cheng2023sammed2d . Figure 4a illustrates the visualization of the attention map and a comparison among BiRD huang2024BiRD , ClinKD (ours), and the ground truth. The grounding of ClinKD is more accurate because MCG-RoPE can improve image-text alignment, and at the same time, Pseudo-KD enables ClinKD to better adapt to medical knowledge.
Case study on VQA-RAD. As shown in Figure 4b, we conducted an experiment on VQA-RAD lau2018dataset . Despite the relatively simple annotations, ClinKD provides fine-grained answers with three accurate groundings, demonstrating its precision in response.
To assess the models’ attention to pathological features, we conducted a comparative analysis of three RoPE variants. The results indicate that 2D-RoPE su2023roformerenhancedtransformerrotary tends to show low attention on pathological regions. In contrast, RoPE-Mixed heo2024rotarypositionembeddingvision demonstrates a greater emphasis on diagonal features. Notably, our Med-CLIP Guided RoPE specifically targets and enhances the model’s attention to pathological regions, thereby facilitating more accurate alignment between text and images, especially in the identification of relevant medical features.
(a) Case study on Med-GRIT-Test30 ye2023samed2d20m ; cheng2023sammed2d .
(b) Case study on VQA-RAD lau2018dataset .
Figure 4: Case Studies on Med-GRIT-Test30 ye2023samed2d20m ; cheng2023sammed2d and VQA-RAD lau2018dataset .
5 Conclusion
In this work, we propose ClinKD, a novel medical distiller centered on Adaptive Confidence-Margin Curriculum Pseudo-KD and Reflective Correction Training. In addition, we propose the Med-CLIP Guided RoPE to improve image-text alignment. They are designed to enhance the performance on fine-grained multi-task datasets in Med-VQA domain. Extensive experiments demonstrate that our Med-CLIP Guided RoPE achieves superior image-text alignment, while the Adaptive Confidence-Margin Curriculum Pseudo-KD effectively bridges prior medical knowledge gaps. These mechanisms synergistically enhance the model’s medical knowledge adaptation capabilities.
In future work, further refinements to the distillation process, along with the incorporation of more comprehensive medical knowledge and higher-quality datasets, could significantly improve model’s robustness and reduce the model’s reliance on manually labeled data.